George drives the first 30 miles of a trip at a constant rate of 40 miles per hour. If he drives the remaining 75 miles of the trip at a constant rate of 50 miles per hour, what is his average speed for the entire trip?
Answer by mananth(16949)

30 miles at 40 mph
time taken at this speed = 30/40 = 3/4 hours
..
75 miles at 50mph
time taken = 75/50 = 3/2 hours
Total time taken = 3/4 +3/2 hours=3+6 /4
=9/4 hours
Total distance traveled = 30+75 = 105 miles
Average speed = 105/ 9/4
=105*4/9
46.6 mph is the average speed